In one year, Steve earned the same interest from an investment at 5% annual interest as an investment at 7% annual interest. He invested $1200 less in the 7% account. How much money was invested in each account?
Answer by checkley77(12844)

.05(X+1200)=.07X
.05X+60=.07X
.05X-.07X=-60
-.02X=-60
X=-60/-.02
X=$3,000 ANS. FOR THE AMOUNT INVESTED @7%
3000+1200=4200 ANS. FOR THE AMOUNT INVESTED @ 5%.
PROOF:
.05*4200=.07*3000
210=210